OpenStudy (anonymous):

Find the absolute extreme values of each function on the interval. y=10-8x^2on [-2,4]

geerky42 (geerky42):

Absolute extreme values is where f'(x) = 0. So take derivative of 10-8x².\[ f'(x) = \dfrac{d}{dx}10-8x^2 = -16x\]\[ -16x = 0\]\[x = 0\] \[f(0) = -10\]So the absolute extreme values of this function occurs at (0, -10)